# Inference for multiple treatment effects using confounder importance learning

### v4

@@ Abstract.
We address modelling and computational issues for multiple treatment effect inference under many potential confounders. A primary issue relates to preventing harmful effects from omitting relevant covariates (under-selection), while not running into over-selection issues that introduce substantial variance and a bias related to the non-random over-inclusion of covariates. We propose a novel empirical Bayes framework for Bayesian model averaging that learns from data the extent to which the inclusion of key covariates should be encouraged, specifically those highly associated to the treatments. A key challenge is computational. We develop fast algorithms, including an Expectation-Propagation variational approximation and simple stochastic gradient optimization algorithms, to learn the hyper-parameters from data. Our framework uses widely-used ingredients and largely existing software, and it is implemented within the R package mombf featured on CRAN. This work is motivated by and is illustrated in two applications. The first is the association between salary variation and discriminatory factors. The second, that has been debated in previous works, is the association between abortion policies and crime. Our approach provides insights that differ from previous analyses especially in situations with weaker treatment effects.
@@ 1. Introduction
We consider a problem of fundamental importance in applied research, that of evaluating the joint effect, if any, of multiple treatments on a response variable when working with observational data and when there are many potential adjustment covariates. We also showcase two applications. One is the association between salary variation and discriminating factors such as gender, race and country of birth, and how this has evolved over time. Another example relates to a debate on a possible association between planned parenthood (as measured by abortion rates) and crime.
Our upcoming discussion is already relevant when the goal is to simply study conditional associations between outcome and treatments (given the covariates). It however becomes critical for causal inference, whereby relying on the so-called non-interference and no unmeasured confounding assumptions, one may identify causal treatment effects from a regression model, provided one selects the necessary covariates (see Antonelli and Dominici (2021) for a recent review). Following standard terminology, we refer by confounders to covariates that are directly associated with both treatment(s) and the response, and by instruments to covariates that correlate with the treatment(s) but are conditionally independent of the outcome, given the treatment(s). Briefly, inferring treatment effects using observational data is subject to selection biases and unmeasured confounding, hence it is important to consider a large number of covariates and to select a suitable subset thereof. It is also important to allow for interaction terms and other transformations, such as polynomial or spline terms to obtain a flexible model and account for possible heterogeneous effects (Caron et al., 2022). As a result, a popular approach to treatment effect inference is to use high-dimensional regression techniques, and we pursue this direction in this article. To put matters in perspective, our analysis of salary variation involves four treatments and hundreds of covariates, and we consider treatment effect heterogeneity across states, whereas for the crime data we have one treatment and 295 covariates.
We model the dependence of the response $y_{i}\sim p(y_{i};\eta_{i},\phi)$ on $t=1,\ldots,T$ treatments $d_{i,t}$ and $j=1,\ldots,J$ covariates $x_{i,j}$, via
@@
where $p(y_{i};\eta_{i},\phi)$ defines a generalized linear model with linear predictor $\eta_{i}$ and dispersion parameter $\phi$ (i.e. the error variance in the Gaussian case, and a known $\phi=1$ in logistic and Poisson regression models). Whereas from an interpretational and policy making point of view the distinction between treatments and covariates is clear, statistically the difference is one of priorities: we are primarily interested in inference for the former, i.e the set of $\alpha_{t}$’s in (1), including uncertainty quantification such as high probability intervals for $\alpha_{t}$, whereas the latter are considered for potential inclusion to avoid omitted variable biases and to allow for flexible regression functions. Although our primary interest is in average treatment effects, it is possible to consider heterogeneous treatment effects by incorporating into $d_{i,t}$ interactions between treatments and covariates. In our salary example we illustrate this by considering interactions between the four primary treatments and state. Importantly, such interactions are added with an add-to-zero constraint. This ensures that the $\alpha_{t}$ associated to a primary treatment quantifies its corresponding average treatment effect, whereas the $\alpha_{t}$’s associated to interactions quantify deviations from the average treatment effect.
@@
Our main interest is in scenarios where the number of covariates $J$ is large. This setting spurred significant interest due to the observation that standard shrinkage and selection methods for learning (1), such as LASSO and Bayesian Model Averaging (BMA), can have an undesirable behavior for treatment effect inference. Intuitively, when some confounders are strongly associated with the treatments, standard high-dimensional methods may fail to include said confounders (or even the treatments) in (1), resulting in significant omitted variable biases in estimating the treatment effects. Two seminal works where these issues are discussed and addressed are Belloni et al. (2014) and Wang et al. (2012). Both works arguably set the basis for subsequent literature, and both consider a single treatment setting ($T=1$). Belloni et al. (2014) proposed a double-LASSO (DL) approach where one regresses separately the response and the treatment on the covariates via the LASSO, takes the covariates with a non-zero estimated effect either on the treatment or the response, and in a second step fits a model like (1) by OLS with these selected covariates. Notably, this treatment effect estimator is asymptotically normally-distributed and has a variance that can be estimated from data. In a similar spirit Wang et al. (2012) proposed Bayesian adjustment for confounders (BAC), which models jointly the response and treatments and uses a prior distribution that encourages covariates to be simultaneously selected in the two regression models.
@@
The main idea in DL, BAC and much of the subsequent literature (reviewed below) is that, by including covariates that are associated to the treatment, one ameliorates omitted variable biases. A key issue motivating our work is that, by protecting oneself against covariate under-selection, one may force (or encourage) the inclusion of instruments, i.e. covariates for which $\beta_{j}=0$ in (1). Under such covariate over-selection treatment effects remain identifiable, however it leads to a problematic variance inflation, see De Luna et al. (2011); Lefebvre et al. (2014); Talbot et al. (2015); Henckel et al. (2022). Adding instruments can severely inflate the treatment effect mean squared error (MSE), and reduce the power to detect weaker effects. To gain intuition, consider a setting with a fixed number of covariates $J$. There, a classical strategy is to fit one model including all covariates to obtain unbiased treatment effects, potentially at the cost of high variance. Specifically, the variance inflation factor for a least-squares estimator of $\alpha_{t}$ is given by $(1-R_{t}^{2})^{-1}$, where $R_{t}$ is the multiple $R^{2}$ coefficient for regressing treatment $t$ on the covariates. Belloni et al. (2014) explain that, when $J$ is fixed, their approach is asymptotically first-order equivalent to fitting a model with all covariates, and hence incurs the mentioned variance inflation factor. A second issue is a more subtle, and to the best of our knowledge yet unnoticed, over-selection bias. Namely, that the inclusion of covariates in (1) which are screened out to be correlated with the treatments and the response may lead to biased inference for the treatment effects. In our experience said over-selection bias is not a major issue in practice, further the results in Belloni et al. (2014) prove that it vanishes asymptotically, hence we defer further discussion to Section D.
@@
Related literature includes Farrell (2015), who adapted the DL framework by using a robust estimator to safeguard from mistakes in the double selection step, and Shortreed and Ertefaie (2017), who employed a two-step adaptive LASSO approach. Chernozhukov et al. (2018) extended DL by introducing a de-biasing step, and cross-fitting to ameliorate false positive inclusion of covariates. As an alternative to these two-step approaches, Ertefaie et al. (2015) used a joint likelihood $L_{1}$ penalization on the outcome and treatment regressions. On the Bayesian side, Lefebvre et al. (2014) discussed how to set the BAC hyper-parameter $\omega$ in a data-based manner to improve the treatment effect MSE. When $\omega=\infty$, the outcome equation includes any covariate associated with the treatment, which akin to DL reduces omitted variable bias at the cost of potentially inflating MSE. The authors warn that the results are sensitive to using 1/2 of the data in their sample-splitting strategy, and of computational challenges if one wanted to consider $T>1$ treatments, further they use a leaps-and-bounds model search that only accommodates up to 31 covariates. Wang et al. (2015) extended BAC to generalized linear models and considered pairwise interactions between the treatment and covariates. The authors used the same prior as BAC and focused on the hyper-parameter choice $\omega=\infty$, which as discussed can be problematic. Talbot et al. (2015) propose a similar framework to BAC where prior probabilities deter the inclusion of instruments to reduce the inclusion of instruments, however said prior probabilities still require a tuning hyper-parameter playing a role similar to $\omega$ in BAC. A proposal related to ours is the ACPME method of Wilson et al. (2018). The framework considers $T>1$ treatments and the prior inclusion probability for covariate $j$ depends, via logistic regression, on a measure of dependence between $j$ and the treatments. Said measure is given by regressing the controls on the treatments and taking the corresponding residual covariance matrix. Analogously to BAC, prior inclusion probabilities are controlled by a tuning parameter, which by default it set such that the average penalty for covariate inclusion is the same as that for the Bayesian information criterion. Antonelli et al. (2019) proposed continuous spike-and-slab Laplace priors on high-dimensional covariates. The framework is designed to reduce the shrinkage to zero of the coefficients of covariates that are associated to the treatment. They discuss how to elicit hyper-parameter values to help shrink the effects of instruments. In a different thread, Hahn et al. (2018) proposed a shrinkage prior framework based on re-parameterizing a joint outcome and treatment regression, designed to improve estimation biases, and Hahn et al. (2020) considered Bayesian causal forests. Overall, a recurrent issue in the Bayesian literature is how to set hyper-parameter values such that one includes all confounders and hence identifies the treatment effects, while preventing variance inflation due to selecting instruments.
@@
Our main contributions are proposing a novel Bayesian model, and an associated computational framework, to learn hyper-parameters that determine whether and how prior covariate inclusion probabilities should relate to their association with the treatments. Our framework, to which we refer by Confounder Importance Learning (CIL), is designed to deal with both over- and under-selection, in both high and low confounding situations. Figure 1 is a first illustration of its merits (see Section 4.1 for details). As discussed in Belloni et al. (2014) and Wang et al. (2012), due to under-selection bias standard LASSO and BMA suffer from high MSE in high-confounding settings, whereas their proposed DL and BAC (respectively) attain much lower MSE. In low-confounding settings however the reverse is true, here DL and BAC have high MSE due to over-selection variance. CIL attains low MSE across the weak-to-strong confounding spectrum.ME CIL can also consider multiple treatments, a setting that has received less attention in the literature. Our model has similarities to that in ACPME, the key difference being that we learn hyper-parameters from the data (and subtler differences discussed in Section 2). However, in Figure 1 we observe that the two methods show different result patterns in the single treatment case, as ACPME closely mimics the behavior of BAC across the board. Relative to Lefebvre et al. (2014), we learn hyper-parameters using the marginal likelihood associated to a Bayesian model rather than a training-test data split. A key ingredient relative to earlier proposals are two scalable computational algorithms, one based on MCMC and the other on a variational approximation. In principle evaluating the (log) marginal likelihood requires a costly sum over $2^{T+J}$ models. We show that, under our proposed prior model, its gradient only requires a sum over $J$ terms that involves only marginal posterior inclusion probabilities. We further propose an expectation-propagation (EP) approximation that bypasses the need to re-estimate said marginal posterior probabilities, which would typically require MCMC. For example, BAC and ACPME failed to return a solution in our salary example after 2 days, whereas our CIL could complete the task in 8 hours and 33 minutes on the largest dataset of 2010. CIL can be easily implemented using existing software, and we provide an implementation within the R package mombf (Rossell et al., 2023).
@@
Figure 1. Parameter root MSE relative to an oracle OLS, for a single treatment, considering strong ($\alpha=1$), weak $(\alpha=1/3)$ and no effect $(\alpha=0)$. In all panels, $n=100$, $J=49$ and the response and treatment are simulated from a linear regression model based on 6 active covariates each. The $x$-axis is the overlap between the two sets of active covariates varies from 0 (no confounding) to 6 (full confounding). DL is double LASSO, BMA is Bayesian model averaging, BAC is Bayesian Adjustment for Confounding and CIL is Confounder Importance Learning
The remainder of paper is structured as follows. Section 2 details our proposed approach, a Bayesian model averaging where prior inclusion probabilities vary across covariates. Section 3 describes our computational methods. Section 4 shows a simulation study illustrating that standard high-dimensional methods have high MSE in high-confounding settings, whereas other methods designed for treatment effect inference have high MSE in low-confounding settings, and that CIL attains a good MSE throughout. We also apply the methodology to the salary and crime case studies. We illustrate how considering multiple treatments allows to portray, via posterior predictive inference, a measure of joint salary variation due to potentially discriminatory factors. All proofs and additional empirical results are provided as a supplement.
2. Modelling Framework
We model the dependence of the response $y_{i}$ on treatments $\mathbf{d}_{i}=(d_{i,1},\ldots,d_{i,T})$ and covariates $\mathbf{x}_{i}=(x_{i,1},\dots,x_{i,J})$, according to (1). We are primarily interested in inference for the treatment effects ${\bm{\alpha}}=(\alpha_{1},\dots,\alpha_{T})$. We adopt a Bayesian framework where we introduce variable inclusion indicators $\gamma_{j}=\text{I}(\beta_{j}\neq 0)$ and $\delta_{t}=\text{I}(\alpha_{t}\neq 0)$, and define a prior

### v6

@@ Abstract.
We address modelling and computational issues for multiple treatment effect inference under many potential confounders. Our main contribution is providing a trade-off between preventing the omission of relevant confounders, while not running into an over-selection of instruments that significantly inflates variance. We propose a novel empirical Bayes framework for Bayesian model averaging that learns from data the prior inclusion probabilities of key covariates. Our framework sets a data-dependent prior that asymptotically matches the true amount of confounding in the data, as measured by a novel confounding coefficient. A key challenge is computational. We develop fast algorithms, using an exact gradient of the marginal likelihood that has linear cost in the number of covariates, and a variational counterpart. Our framework uses widely-used ingredients and largely existing software, and it is implemented within the R package mombf. We illustrate our work with two applications. The first is the association between salary variation and discriminatory factors. The second, that has been debated in previous works, is the association between abortion policies and crime. Our approach provides insights that differ from previous analyses especially in situations with weaker treatment effects.
@@ 1. Introduction
We consider a fundamental problem in applied research, that of evaluating the joint association, if any, of multiple treatments on an outcome when working with observational data and when there are many potential adjustment covariates. In such settings, it is common to use generalized linear models (GLMs) or additive models. Our discussion also applies to causal inference, whereby relying on the so-called non-interference and no unmeasured confounding assumptions, one may identify causal treatment effects from a regression model, provided one selects the necessary covariates and models properly their association with the outcome (see Antonelli and Dominici (2021) for a review). Following standard terminology, we refer by confounders to covariates that are truly associated with both treatment(s) and the response (given other covariates), and by instruments to covariates that correlate with the treatment(s) but are conditionally independent of the outcome. A common approach to estimate conditional associations is to learn which covariates are confounders using high-dimensional regression, and we pursue this direction in this article. As we shall discuss, a key issue that we address is attaining a good trade-off between avoiding omitted variable biases and variance inflation driven by instruments. We achieve this by setting a data-dependent prior via empirical Bayes, and proposing highly efficient algorithms to estimate the required hyper-parameters. We show that asymptotically the hyper-parameter estimates adapt to the true sparsity in the data, and capture a novel measure of confounding that we introduce in this article.
We model the dependence of the outcome $y_{i}\sim p(y_{i};\eta_{i},\phi)$ on $t=1,\ldots,T$ treatments $d_{i,t}$ and $j=1,\ldots,J$ covariates $x_{i,j}$, via
@@
where $p(y_{i};\eta_{i},\phi)$ defines a GLM with linear predictor $\eta_{i}$ and dispersion parameter $\phi$ (i.e. the error variance in the Gaussian case, and a known $\phi=1$ in logistic and Poisson regression). Whereas from an interpretational and policy making point of view the distinction between treatments and covariates is clear, statistically the difference is one of priorities: we are primarily interested in inference for treatment effects ($\alpha_{t}$’s in (1)), including uncertainty quantification, whereas the $\beta_{j}$’s are considered to avoid omitted variable biases and to allow for flexible regression functions. Although our primary interest is in average treatment effects, it is possible to consider heterogeneous effects by incorporating into $d_{i,t}$ interactions between treatments and covariates. In our salary example we illustrate this by considering interactions between the four primary treatments and state. Importantly, such interactions are added with an add-to-zero constraint. This ensures that the $\alpha_{t}$ associated to a primary treatment quantifies its corresponding average treatment effect, whereas the $\alpha_{t}$’s associated to interactions quantify deviations from the average treatment effect.
@@
Our main interest is in scenarios where the number of covariates $J$ is large. This setting spurred significant interest due to the observation that standard shrinkage and selection methods for learning (1), such as LASSO and Bayesian Model Averaging (BMA), can have an undesirable behavior for treatment effect inference. When many confounders are strongly associated with the treatments, a situation that we refer to as high-confounding, standard high-dimensional methods may fail to include said confounders (or even the treatments) in (1), resulting in significant omitted variable biases. Two seminal works are Belloni et al. (2014) and Wang et al. (2012). Both set the basis for subsequent literature, and both consider a single treatment setting ($T=1$). Belloni et al. (2014) proposed a double-LASSO (DL) approach where one regresses separately the outcome and the treatment on the covariates via the LASSO, takes the covariates with a non-zero estimated effect either on the treatment or the outcome, and in a second step fits a model like (1) by maximum likelihood estimation (MLE) with these selected covariates. Notably, this treatment effect estimator is asymptotically normal and has a variance that can be estimated from data. In a similar spirit Wang et al. (2012) proposed Bayesian adjustment for confounders (BAC), which models jointly the outcome and treatments and uses a prior distribution that encourages covariates to be simultaneously selected in the two regression models.
The main idea in DL, BAC and subsequent literature (reviewed below) is that, by including covariates that are associated to the treatment, one ameliorates omitted variable biases. A related notion called regularization-induced confounding (RIC) refers to estimation biases due to not properly accounting for confounders, due to the prior over-shrinking in specific directions (Hahn et al., 2018; Linero and Antonelli, 2023). This notion is related to the omitted variable bias discussed by Belloni et al. (2014), i.e. the concern is not properly handling the confounders, and is addressed by linking the outcome and treatment models.
@@
A key distinction motivating our work is that, by protecting oneself against omitted variables, one may force (or encourage) the inclusion of instruments, i.e. covariates for which truly $\beta_{j}=0$ in (1). Under such covariate over-selection treatment effects remain identifiable, however there is a problematic variance inflation, see De Luna et al. (2011); Lefebvre et al. (2014); Zigler and Dominici (2014); Talbot et al. (2015); Henckel et al. (2022). We argue that one should try to reach a compromise between handling properly the confounders and the instruments. Adding instruments can severely inflate the treatment effect mean squared error (MSE), and reduce the power to detect weaker effects. To gain intuition, consider a setting with a fixed number of covariates $J$. A classical strategy is to fit one model including all covariates to obtain unbiased treatment effects, potentially at the cost of high variance. Specifically, the variance inflation factor for a least-squares estimator of $\alpha_{t}$ is given by $(1-R_{t}^{2})^{-1}$, where $R_{t}$ is the multiple $R^{2}$ coefficient for regressing treatment $t$ on the covariates. Hence, if one has instruments that accurately predict the treatment, $R_{t}^{2}$ is close to 1 and variance inflation is severe. Belloni et al. (2014) explain that, when $J$ is fixed, their approach is asymptotically first-order equivalent to fitting a model with all covariates, and hence incurs variance inflation. A second issue is a more subtle over-selection bias that received less attention (but see Zigler and Dominici (2014) for a brief mention). Namely, including covariates in (1) that are correlated with the treatments and the outcome may lead to biased inference. In our experience over-selection bias is not a major issue in practice, further the results in Belloni et al. (2014) prove that it vanishes asymptotically, hence we defer further discussion to Section S1.
@@
Related literature includes Farrell (2015), who adapted the DL framework by using a robust estimator to safeguard from mistakes in the double selection step, and Shortreed and Ertefaie (2017), who employed a two-step adaptive LASSO approach. Chernozhukov et al. (2018) extended DL by introducing a de-biasing step, and cross-fitting to ameliorate false positive inclusion of covariates. On the Bayesian side, Lefebvre et al. (2014) discussed how to set the BAC hyper-parameter $\omega$ in a data-based manner to improve the treatment effect MSE. When $\omega=\infty$, the outcome equation includes any covariate associated with the treatment, which akin to DL reduces omitted variable bias at the cost of potential variance inflation. The authors warn that the results are sensitive to using half of the data in their sample-splitting strategy, and of computational challenges if one wanted to consider $T>1$ treatments, further they use a leaps-and-bounds model search that only accommodates up to 31 covariates. Wang et al. (2015) extended BAC to GLMs and considered pairwise interactions between the treatment and covariates. The authors used the same prior as BAC and focused on the hyper-parameter choice $\omega=\infty$, which as discussed can be problematic. Talbot et al. (2015) propose a similar framework to BAC where prior probabilities deter the inclusion of instruments to reduce the inclusion of instruments, however said prior probabilities still require a tuning hyper-parameter playing a role similar to $\omega$ in BAC. A proposal that is closest to ours is the ACPME method of Wilson et al. (2018). The framework considers $T>1$ treatments and the prior inclusion probability for covariate $j$ depends, via logistic regression, on a measure of dependence between $j$ and the treatments. Analogously to BAC, prior inclusion probabilities are controlled by a tuning parameter, which by default sets the same average penalty for covariate inclusion as the Bayesian information criterion. A key difference is that Wilson et al. (2018) do not use the outcome data to drive prior inclusion probabilities. They assume that any control associated to the treatment(s) is likely to be needed in the outcome equation, to an extent driven by a user-defined hyper-parameter. In low-confounding settings where there are many instruments, this assumption is violated. Instead, we use the outcome to set data-dependent prior inclusion probabilities, by learning whether there truly is high or low confounding (hence the naming Confounding Importance Learning). See Section 2.1 for further comparisons with ACPME. Antonelli et al. (2019) proposed continuous spike-and-slab Laplace priors on high-dimensional covariates. The framework is designed to reduce the shrinkage to zero for covariates that are associated to the treatment. They discuss how to elicit hyper-parameters to help shrink the effects of instruments. In a different thread, Hahn et al. (2018) proposed shrinkage priors based on re-parameterizing a joint outcome and treatment regression.
@@
Overall, a recurrent issue is how to set hyper-parameters to avoid omitted-variable biases but also prevent variance inflation due to selecting instruments. Our main contribution is a novel framework that sets (data-dependent) prior inclusion probabilities to balance these two competing goals. We prove that our framework sets prior inclusion probabilities which, using empirical Bayes, asymptotically reflect a novel confounding coefficient introduced here. Said coefficient reflects whether one is in a situation with many confounders (high confounding), many instruments (low confounding), or neither (neutral confounding). Our framework, which we call Confounder Importance Learning (CIL), is designed to deal with both over- and under-selection, in both high and low confounding situations. Figure 1 is a first illustration of its merits (see Section 5.1 for details). As discussed, due to omitted-variable bias standard LASSO and BMA suffer from high MSE in high-confounding settings, whereas DL and BAC attain much lower MSE. In low-confounding settings however the reverse is true, here DL and BAC have high MSE due to over-selection variance. CIL attains low MSE across the high-to-low confounding spectrum. CIL can also consider multiple treatments, a setting that has received less attention in the literature. Although our model has similarities to ACPME, Figure 1 shows that the two methods behave quite differently, as ACPME closely mimics the behavior of BAC. Relative to Lefebvre et al. (2014), we learn hyper-parameters using the marginal likelihood associated to a Bayesian model rather than a training-test data split, which is integral in showing that our prior probabilities asymptotically match the (unknown) true confounding coefficient. Another important contribution are two scalable computational algorithms, based on MCMC and on a variational approximation. In principle evaluating the (log) marginal likelihood requires a costly sum over $2^{T+J}$ models. We show that, under our proposed prior model, its gradient only requires a sum over $J$ terms that involves only marginal posterior inclusion probabilities. We further propose an expectation-propagation (EP) approximation that bypasses the need to re-estimate said marginal posterior probabilities, which would typically require MCMC. For example, BAC and ACPME failed to return a solution in our salary example after 2 days, whereas our CIL could complete the task in 8 hours and 33 minutes on the largest dataset of 2010. CIL can be easily implemented using existing software, and we provide an implementation within the R package mombf (Rossell et al., 2023).
@@
Figure 1. Parameter root MSE relative to an oracle OLS, for a single treatment, considering strong ($\alpha=1$), weak $(\alpha=1/3)$ and no effect $(\alpha=0)$. In all panels, $n=100$, $J=49$ and the outcome and treatment are simulated from a linear regression model based on 6 active covariates each. The $x$-axis is the overlap between the two sets of active covariates varies from 0 (no confounding) to 6 (full confounding). DL is double LASSO, BMA is Bayesian model averaging, BAC is Bayesian Adjustment for Confounding and CIL is Confounder Importance Learning
This paper is structured as follows. Section 2 details our proposed approach, a Bayesian model averaging where prior inclusion probabilities vary across covariates. It also introduces a confounding coefficient that plays a pivotal role in interpreting our methodology, and how it differs from current literature. Section 3 describes our computational methods, and how empirical Bayes seeks to match the prior mean of the confounding coefficient to its posterior mean. Section 4 shows that the latter converges to the data-generating confounding coefficient, allowing for model misspecification, in finite-dimensional settings. Section 5 shows simulations, and a salary and a crime case study. All proofs and additional empirical results are provided as a supplement.
2. Framework
2.1. Model
We model the dependence of the outcome $y_{i}$ on treatments $\mathbf{d}_{i}=(d_{i,1},\ldots,d_{i,T})$ and covariates $\mathbf{x}_{i}=(x_{i,1},\dots,x_{i,J})$, according to (1). We are primarily interested in inference for the treatment effects ${\bm{\alpha}}=(\alpha_{1},\dots,\alpha_{T})$. We adopt a Bayesian framework where we introduce variable inclusion indicators $\gamma_{j}=\text{I}(\beta_{j}\neq 0)$ and $\delta_{t}=\text{I}(\alpha_{t}\neq 0)$, and define a prior